OpenStudy (anonymous):

find the product (-4)*3*(-2)*6 with steps

OpenStudy (anonymous):

(-4)(3)(-2)(6) =(-12)(-2)(6) =(24)(6) =144

OpenStudy (anonymous):

Multiply 3 by 6 to get 18. 18*-4*-2 18*8 answer: 144

OpenStudy (mayankdevnani):

\[(-4)\times3\times(-2)\times6\] \[(-12)\times(-12)\] \[(-12)^{2}\] \[144\]

OpenStudy (mayankdevnani):

@mano345